what are the horizontal intercepts of the graph of the quadratic function
32=x^2-8x+12

Answer by stanbon(75887)

what are the horizontal intercepts of the graph of the quadratic function
32=x^2-8x+12
------
x^2 -8x -20 = 0
------
Factor:
(x-10)(x+2) = 0
x = 10 or x = -2
--------------------------
Cheers,
Stan H.
--------------------------